OpenStudy (anonymous):

Given a cube with a volume of 27 cm3, what is the volume of a square pyramid that can fit perfectly inside the cube? (4 points) 3cm3 6cm3 9cm3 12cm3 I KNOW that the answer is 9cm3, my issue is that I don't know how. I always thought that a the volume of a pyramid was 1/6(b^3) or the volume of a cube but apparently that's not right. I'm confused on how to reach the answer if someone could please show and explain to me the process of getting to 9cm^3

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

@nciezki12 I know, how how did you reach that conclusion?

OpenStudy (anonymous):

27 divided by 3 = 9

OpenStudy (anonymous):

Yes why did you choose to multiply it by 1/3 though?

OpenStudy (anonymous):

Is the volume of a pyramid = to 1/3 of the volume of a cube? If so why?

OpenStudy (anonymous):

DOD this is yours qustion not mine

OpenStudy (anonymous):

DOD?

OpenStudy (anonymous):

@nciezki12 what's DOD?

OpenStudy (anonymous):

this is you q

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

@Mrs.Peter

OpenStudy (anonymous):

u ther

OpenStudy (anonymous):

@nciezki12 Yes. :)

OpenStudy (anonymous):

this is your question

OpenStudy (anonymous):

@iBacon @iGreen @iBacon @iGreen

OpenStudy (anonymous):

How do you know that is is C?

OpenStudy (anonymous):

he is telling me how did you get to that conclusion

OpenStudy (anonymous):

i said this is his qustion

OpenStudy (anonymous):

you are supposed so explain to him how to get the answer.

OpenStudy (anonymous):

and i told him 27 divided by 3 = 9 that is my exlanain

OpenStudy (anonymous):

1/3(B)(h) is the formula for the volume of a pyramid. The volume of a square is b^3 and we know that because the volume of the cube is 27 that b = 3. B = side times side so 1/3 (3*3)(h) = 1/3(9) in which h = 3 so 1/3 27 the answer is 9

OpenStudy (anonymous):

I get it.

OpenStudy (anonymous):

Sorry, I had to work it out xD I get it now. Im so sorry

OpenStudy (anonymous):

@Mrs.Peter V=lwh/3 :)

OpenStudy (anonymous):

Will this always work ? 1/3 the volume of a cube equals the volume of a pyramid that is inscribed within the cube?

OpenStudy (anonymous):

yes it will

OpenStudy (anonymous):

Just multiply LxWxH /3

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@nciezki12 @iBacon thank you guys so much for your help! I was very confused but not anymore. :)

OpenStudy (anonymous):

But it gave you the LxWxH. So just divide it by 3. :)

OpenStudy (anonymous):

@nciezki12 next time help him don't just tell him the answer it's against CoC

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@nciezki12 @iBacon Thanks again, I thought both of you did a great job in helping me :)

OpenStudy (anonymous):

No problem. If you need help on another question tag me in it. :)

OpenStudy (anonymous):

@iBacon @Mrs.Peter can u help me

OpenStudy (anonymous):

pls

OpenStudy (anonymous):

Yes, of course

OpenStudy (anonymous):

start a new question don't comment on his

OpenStudy (anonymous):

@nciezki12 link me the question?

OpenStudy (anonymous):

@Mrs.Peter close question :)

OpenStudy (anonymous):

kk, thanks again you guys God bless : - )

OpenStudy (anonymous):

i did it